User Guide for VVencode/VVdecode 1

1 Guide to using VVcode

• general introduction

1.1 VVencode command options

• described in a generic manner with operating system specifics at the end of the section The general syntax of the VVencode command line is:

vvencode [options] input-file [options] [output-file] [options]

There are two parameters and many optional qualifiers. The parameters are:

input-file
output-file

The command qualifiers modify the behaviour of VVcode and are all optional. Qualifiers may appear before, between or after the parameters. The qualifiers acceptable to VVencode are: debug Write copious debugging output to the standard error device or to the log file if the

+log qualifier was specified. This qualifier would normally only be used when trying to report a VVencode problem and may indeed have been disabled for the version that you are using.

encoding_table file-spec

Read the encoding table to be used from the file ‘file-spec’. The file must contain the 64 characters to be used by VVencode. If this qualifier is not specified, the default VVencode encoding table will be used.

format file-format

Explicitly specify the file format that should be assumed by VVencode when it is reading the input file. The file-format value must be one of ‘fixed’, ‘stream’ and ‘variable’. This qualifier would only be used if it was necessary to override the value determined automatically by VVencode.

header_filespec file-spec

help VVencode will display a usage message on the standard error device and exit.

log file-spec
mode file-mode

Explicitly specify the file mode that should be assumed by VVencode when it is reading the input file. The file-format value must be one of ‘binary’ and ‘text’. This qualifier would only be used if it was necessary to override the value determined automatically by VVencode.

overwrite
record_length
split_size

timestamp time
translation_file file-spec

1.2 VVdecode command options

The general syntax of the VVdecode command line is:

     vvdecode [options] input-file [options] [output-file] [options]

1.3 Operating system specific considerations

1.3.1 MS-DOS

• qualifiers introduced with ‘/’ or ‘-’ and must be separated from each other by whitespace • values separated from qualifier by ‘=’ or ‘:’

• optional values supported

• VVcode will act as a filter when ‘-’ is specified as the input or output file

• optional file specification for ‘log’ qualifier

• examples of VVencode usage

1.3.2 OS/2

• qualifiers introduced with ‘/’ or ‘-’ and must be separated from each other by whitespace • values separated from qualifier by ‘=’ or ‘:’

• optional values supported

• VVcode will act as a filter when ‘-’ is specified as the input or output file

• optional file specification for ‘log’ qualifier

• GCC 2.1 implementation does not support file timestamping

• HPFS file names supported

• OS/2 extended attributes not supported yet

• examples of VVencode usage

1.3.3 Unix

• qualifiers introduced with ‘-’ or ‘+’ and must be separated from each other by whitespace • values separated from qualifier by whitespace

• optional values not supported

• VVcode will act as a filter when ‘-’ is specified as the input or output file

• examples of VVencode usage

1.3.4 VAX/VMS

• qualifiers introduced with ‘/’

• values separated from qualifier by ‘=’

• standard VMS CLD conventions

• optional values supported

• VVcode will act as a filter when ‘-’ is specified as the input or output file though meaningless on VMS

• optional file specification for ‘log’ qualifier

• examples of VVencode usage

1.3.5 VM/CMS

• qualifiers introduced with ‘(’ and must be separated from each other by whitespace

• values separated from qualifier by ‘=’ or ‘:’

• command line driven through REXX EXEC

• optional values supported

• VVcode does not support input or output from/to standard I/O

• optional file specification for ‘log’ qualifier

• examples of VVencode usage

1.4 Transferring files between systems

• file type and format

• character set

• electronic mail

• timestamping

1.5 Unusual applications of VVcode

• character set conversion

• file type conversion

• VMS file format conversion vs. CONVERT

Table of Contents

1 Guide to using VVcode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1.1 VVencode command options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 VVdecode command options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Operating system specific considerations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.3.1 MS-DOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3.2 OS/2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3.3 Unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3.4 VAX/VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.5 VM/CMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.4 Transferring files between systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.5 Unusual applications of VVcode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4